How do I solve these questions? I know how I have to use the geometric sequence, but I'm stumped. The question is, "True or False, are these equal?"
1000
Σ 7 (3/4)^n
n=2

998
63/16 Σ (3/4)^n
n=0
(I cannot get my computer to edit this question correctly. Just to clarify: The 63/16 is in front of the sigma sign. The 998 is above sigma and n=0 below. The (3/4)^n is to the right)
THANK YOU, THANK YOU, THANK YOU!!

Answer by Emsley(35)

First series 1st term=63/16 Common ratio=3/4 No of terms=999
So sum=63/16(1-3/4^999)/1-3/4=63/4(1-3/4^999)
Second series 1st term=1 Common ratio=3/4 No of terms=999

sum=63S/16*1(1-3/4^999)/(1-3/4)=63/16(1-3/4^999) So TRUE